A salad bars customers are charged for the amount of salad they take. Sampling suggests that the amount of salad taken is uniformly distributed between 5 ounces and 15 ounces. What is the probability that a customer will take between 12 and 15 ounces of salad?


1
Expert's answer
2022-05-30T16:22:36-0400
"P(12\le X\le 15)=\dfrac{15-12}{15-5}=0.3"